What does 2 log2 8 - log4 256 mean or equals becuase I have no idea on what it means?...Help Me please
Answer by stanbon(48535)

What does 2 log2 8 - log4 256 mean or equals becuase I have no idea on what it means?...
--------------
One step at a time:
log means "exponent" or "power"
----------------
2log2 8 means "twice the exponent of 2 that gives you the number 8.
----------------
So what exponent of 2 gives you the number 8?
2^3 = 8
So log2 8 is 3
Twice that is 6
---------------------
How about log4 256
4^4 = 246
So log4 256 is 4
---------------------
So 2log2 8 - log4 256 means 2*3-4 = 2
=======================================
Cheers,
Stan H.